1) A junior staff member has provided you with the following balance sheet information for 2014 & 2015 along with income statement information for 2015 2014 $1,000 Catego 2015 $1,200 5,600 Cate 2015 $10,000 Sales Cost of Goods Sold Interest Expense Current Assets 5,000 600 $4,800 $0 Net Fixed Assets Accounts Payable 800 Accrued Expenses Long-term Debt 500 3,000 600 3,300 Depreciation Tax Rate $1,200 30% a) Calculate Net Working Capital for 2015 b)Calculate the Operating Cash Flow for 2015 c) Calculate the Free Cash Flow for 2015 d) Clearly the junior staffer left out a critical element of the balance sheet for 2014 & 2015. What was left out and what must it be equal to in 2014 & 2015?

Please refrain from using Excel thanks!
Answer #1

Part (a)

Net Working Capital, NWC = Current Assets, CA - Current Liabilities, CL

Current Liabilities, CL = Account Payables + Accrued Expenses

NWC2015 = CA2015 - CL2015 = 1,200 - (800 + 600) = - 200

Part (b)

Operating cash flow, OCF = Profits after tax + Depreciation - Increase in working capital

For the year 2015:

Profits after tax = (Sales - Cost of good sold - Depreciation - Interest) x (1 - Tax rate) = (10,000 - 4,800 - 1,200 - 0) x (1 - 30%) = 2,800

Increase in working capital = Working capital of 2015 - working capital of 2014

Working capital of 2015 = - 200 as calculated in part (a)

Working capital of 2014 = Current assets in 2014 - Current liabilities in 2014 = 1,000 - (600 + 500) = -100

Increase in working capital = Working capital of 2015 - working capital of 2014 = -200 - (-100) = -100

Hence, OCF = Profits after tax + Depreciation - Increase in working capital = 2,800 + 1,200 - (-100) = 4,100

Part(c)

Free cash flow, FCF = OCF - Investment in Fixed assets

Net fixed assets of 2015 = Net fixed assets of 2014 + Investment in fixed assets during 2015 - Depreciation during 2015

Hence: 5,600 = 5,000 + Investment in fixed assets during 2015 - 1,200

Hence, Investment in fixed assets during 2015 = 5,600 + 1,200 - 5,000 = 1,800

Part (d)

The junior staff has left out the shareholders' equity on the balance sheet. The same should be the balancing figure in the balance sheet.

Shareholders' equity in 2015 = Total assets - Total liabilities = (1,200 + 5,600) - (800 + 600 + 3,300) = 2,100

Shareholders' equity in 2014 = Total assets - Total liabilities = (1,000 + 5,000) - (600 + 500 + 3,000) = 1,900